5. Problems and Applications Q5 The nation of Textilia does not allow imports of clothing. In its equilibrium without trade, a T-shirt costs $24, and the equilibrium quantity is 4 million T-shirts. One day, after reading Adam Smiths The Wealth of Nations while on vacation, the president decides to open the Textilian market to international trade. The market price of a T-shirt falls to the world price of $16. The number of T-shirts consumed in Textilia rises to 8 million, while the number of T-shirts produced declines to 2 million. Use the blue line (circle symbol) to graph the domestic demand for T-shirts in Textilia. Then use the orange line (square symbol) to graph the domestic supply of T-shirts in Textilia. Next, use the black point (plus symbol) to indicate the domestic equilibrium price and quantity before trade. Finally, use the grey line (star symbol) to indicate the world price. Note: Assume the domestic demand and domestic supply curves are linear. 40 36 Domestic Demand 32 28 24 Domestic Supply 20 2 16 a. Domestic Equilibrium 12 World Price 4Complete the following table by calculating the consumer surplus, producer surplus, and total surplus before and after trade. (Hint: Recall that the area of a triangle is 흘 × base × height.) Before Trade After Trade (Millions of Dollars) (Millions of Dollars) Consumer Surplus Producer Surplus Total Surplus As a result of opening up to trade, consumer surplus producer surplus ,and total surplus
Answer #1

(a) Linear demand equation: P = a - bQ

When P = 24, Q = 4 and when P = 16, Q = 8.

24 = a - 4b.......(1)

16 = a - 8b.......(2)

(1) - (2) yields:

8 = 4b

b = 2

a = 16 + 8b [From (2)] = 16 + (8 x 2) = 16 + 16 = 32

Linear demand equation: P = 32 - 2Q

When Q = 0, P = 32 (Vertical intercept) and when P = 0, Q = 32/2 = 16 (Horizontal intercept).

(b) Linear supply equation: P = c + dQ

When P = 24, Q = 4 and when P = 16, Q = 2.

24 = c + 4d.......(1)

16 = c + 2d.......(2)

(1) - (2) yields:

8 = 2d

d = 4

c = 16 - 2d [From (2)] = 16 - (4 x 2) = 16 - 8 = 8

Linear demand equation: P = 8 + 4Q

When Q = 0, P = 8 (Vertical intercept).

(c) In following graph, demand and supply curves are drawn. Equilibrium is at intersection of demand and supply curves with price 24 and quantity 4.

40 36 Domestic Demand 32 28 Domestic Supply 24 20 16 Domestic Equilibrium 12 8 World Price Price of T-shirts

(d)

Consumer surplus (CS) = Area between demand curve and market price

Producer surplus (PS) = Area between supply curve and market price

Total surplus (TS) = CS + PS

(1) Before trade:

CS ($ million) = (1/2) x (32 - 24) x 4 = 2 x 8 = 16

PS ($ million) = (1/2) x (24 - 8) x 4 = 2 x 16 = 32

TS ($ million) = 16 + 32 = 48

(2) After trade:

CS ($ million) = (1/2) x (32 - 16) x 8 = 4 x 16 = 64

PS ($ million) = (1/2) x (16 - 8) x 2 = 1 x 8 = 8

TS ($ million) = 64 + 8 = 72

(e) As a result of trade, CS increases by $48 million (= 64 - 16), PS decreases by $24 million (= 32 - 8) and TS increases by $24 million (= 72 - 48).